In 2006 the population of mexico was 107.4 million. If Mexico's population continues to grow at an annual rate of 1.43%, then the population in 2008 will be (1.0143) to the power 14 million.
a) Find the predicted population in 2020 to the nearest tenth of a million people.

I have had lots of trouble with one because I do not remember the formula

Answer by rfer(16322)

FP=107.4(1.0143)^2
FP=107.4*1.0288
FP=110.5 million
-----------------
FP=107.4(1.0143)^14
FP=107.4*1.219916
FP=131.02 million
I think it is ti the power of 14 years not 14 million.